OpenStudy (anonymous):

What are the solutions to the equation 2x^2 +7x -15?

OpenStudy (radar):

It will factor quite nicely. Use that method, factor, set the products to 0 and solve.

OpenStudy (radar):

The first term is prime so for a start. (x ------)(2x______) The 3rd term is negative so the signs of the the products of 15 will be different. the middle term is + so the + sign will go with the product that will result in a + value. (x + __) (2x - ___) Now mentally figure what will produce -15 when multiplied and when used with the produce of 2x^2, (2 and 1) combine to produce 7. (x + 5) (2x - 3) Now you say 2x^2 + 7x -15 is an equation so I assume their is a = sign somewhere was it 2x^2 + 7x -15 =0 ?? If so then: (x +5) (x - 3) = 0 Now you can solve.

OpenStudy (anonymous):

thankyou I understand perfectly now

OpenStudy (radar):

Good for you, good luck with your studies.

OpenStudy (anonymous):

can you help me with this one system of equations problem?

OpenStudy (radar):

I might, but I am not right all the time. Lets see what you have with the system, if it is a linear system, I may be of some help.

OpenStudy (anonymous):

Jamie is organizing a coed soccer team for a community soccer league. The team must have from 12 to 15 adults, and the number of women must be greater than half the number of me. First write a system of inequalities to model the problem. Then solve the system of inequalities by graphing

OpenStudy (aum):

Typo in the last line of radar's reply to the previous problem: "(x +5) (x - 3) = 0 Now you can solve." Should be (x+5)(2x-3) = 0.

OpenStudy (radar):

Let w = women, and let m = men 12 =<(w + m)=<15 w > m/2 I think that would be two inequaltities. Now graph each one and where they share area on the graph would be the solutions

OpenStudy (radar):

Thanks @aum, I did a typo. I told ludvic I am not right all the time lol.

OpenStudy (radar):

Could you help with the graph question, I am kind of confused. Maybe it is isn't a system afterall, maybe a single expression would do. I am still thinking about it.

OpenStudy (radar):

\[12\le (w >m/2) + m \le 15\] Would this say what the problem is providing?

OpenStudy (aum):

There are 3 inequalities here: 12 <= w + m or w >= -m + 12 w + m <= 15 or w <= -m + 15 w > m/2 Let w be along the y-axis and m be along the x-axis. Graph the three inequalities. Shade the appropriate regions. Find the common shaded region and that will be the solution space.

OpenStudy (radar):

@ludvic you be thinking too. @aum that seems much simpler than the convoluted expression I came up with.

OpenStudy (radar):

@ludvic am leaving the graphing up to you.

OpenStudy (aum):

http://prntscr.com/44lf1m

OpenStudy (anonymous):

Thankyou so much @aum and @radar

OpenStudy (aum):

You are welcome.

OpenStudy (radar):

Thanks @aum, that was a neat graph. I was wondering what it looked like. You're welcome @ludvic

OpenStudy (radar):

@ludvic on the graph y is the women, and x is the men.